# BEVPlace++: Fast, Robust, and Lightweight LiDAR Global Localization for Unmanned Ground Vehicles

### v2

@@ V-B Evaluation Metrics
We use different metrics to evaluate different tasks. For place recognition, we use the recall rate at top-1 following previous works [30, 68, 14]. For each query, we find its Top-1 match from the database. According to a distance threshold of 5 meters [30], we determine whether the prediction is a true positive (TP), a false negative (FN), or a false positive (FP). The recall rate is defined as the ratio of TP over the actual positives, i.e.
@@
For loop closing with loop closure detection and pose estimation, we use the Precision-Recall (PR) curve, average precision, F1 max score, max recall at 100% precision, mean translation errors, and mean rotation errors. The PR curve relies on both recall rates and precision, where precision is computed as the ratio of true positives (TP) over all predicted positives, i.e.
@@
For complete global localization, we evaluate mean translation errors and mean rotation errors. Following [21], we also compute the localization success rate (SR) under a threshold of $(2m,5^{\circ})$.
TABLE III: Recall at Top-1 on the KITTI dataset.
@@
TABLE IV: Recall at Top-1 on the rotated KITTI dataset.
@@
VI Experiments
In this section, we conduct experiments to evaluate the performance of our method in terms of place recognition, loop closing, and complete global localization. We compare our method with state-of-the-art methods including M2DP [33], CVTNet [18], Logg3D-Net [28], BoW3d [53], and LCDNet [20]. Among these methods, BoW3d and LCDNet can estimate poses, while the other 3 methods can only perform place recognition and loop closure detection.
TABLE V: Generalization Performance of Recall at Top-1 on NCLT and MCD datasets.
@@ VI Experiments
TABLE VI: Recall rates on the benchmark dataset.
@@
Implementation details. For all baseline methods, we reproduce their results using their open-source implementations with default setups [^2] https://github.com/LiHeUA/M2DP,https://github.com/BIT-MJY/CVTNet,https://github.com/YungeCui/BoW3D,https://github.com/csiro-robotics/LoGG3D-Net,https://github.com/robot-learning-freiburg/LCDNet.. For BEVPlace++, we use ResNet34 as the backbone in REM. The ResNet is cropped to retain the first three layers (up to conv3_x), resulting in an output channel number $C$ of 128. The number of rotations $N_{R}$ in REM is empirically set to 8. The point cloud crop range $D$ is set to 40 meters, and the grid size $g$ for BEV image generation is 0.4 meters. Consequently, the BEV image has a size of $200\times 200$. We train BEVPlace++ with the AdamW optimizer for 50 epochs. The learning rate is initially set as 1e-4 and decays by a factor of 2 every 10 epochs. The weight decay is set to 1e-3. The method is trained on an RTX 3090 GPU.
VI-A Place Recognition
Performance on KITTI. We only train the methods on KITTI dataset using the database of sequence “00”, which contains 3000 frames. We apply data augmentation by randomly rotating the point clouds to improve the robustness to view changes. As can be seen in Tab. III, our BEVPlace++ outperforms M2DP, Scan Context, BoW3D, and CVTNet with large margins. Logg3d-Net and LCDNet achieve comparable performance to BEVPlace++. However, they perform much worse than BEVPlace++ on the challenging sequence “08” which has a large amount of challenging reverse loops.
Robustness to view changes. To test the robustness against rotational changes, we randomly rotate all the query and database point clouds around the z-axis with an angle range of $[0,2\pi)$ to simulate view changes. As shown in Tab. IV, our BEVPlace++ maintains the highest recall rates, benefiting from our rotational invariant global descriptor designs. CVTNet, Logg3d-Net, and LCDNet are also robust to rotations to some extent. However, BoW3D’s performance significantly degenerated compared to those without rotation changes.
Generalization ability and Long-term performance. We evaluate the methods on NCLT and MCD datasets using models trained on KITTI. For NCLT, we construct the database with the sequence “2012-01-15” that covers most areas of the campus. We then perform place retrieval using the point clouds of other sequences, including the one collected in 2013 across a year. For MCD, we build the database with the sequence “ntu_day_01” and perform place retrieval using other sequences, including the three night sequences. These two datasets are collected using different types of LiDAR scanners compared to that used in KITTI and their point clouds are sparser. Tab. V shows the recall rates at Top-1 on the two datasets. As can be seen, BEVPlace++ achieves high recalls on NCLT regardless of season changes. The compared methods rather have much lower recall rates. BEVPlace++ consistently outperforms other methods on MCD with day-night changes.
TABLE VII: Average precision and F1 max score of loop closure on KITTI and NCLT.
@@ VI-A Place Recognition
LCDNet [40] 95.2 91.8 12.2 0.21 0.47 0.633 0.342 0.0 0.39 1.20 0.621 0.362 0.0 0.37 1.16 0.684 0.321 0.0 0.37 1.26
BEVPlace++ 0.999 0.984 76.4 0.35 0.57 0.963 0.912 24.9 0.34 1.09 0.969 0.916 34.5 0.36 1.19 0.935 0.859 31.2 0.40 1.17
@@
BEVPlace++ 0.955 0.901 63.4 0.40 1.19 0.957 0.894 45.3 0.40 1.61 0.839 0.733 15.8 0.40 1.10 0.959 0.887 45.5 0.44 1.05
@@
Performance on sparse point maps of the RobotCar and Inhouse dataset. The two datasets provide point clouds downsampled to 4096 points and normalized to range $[-1,1]$. They contain coarse position ground truth. While the compared methods need raw points input or accurate pose supervision, we do not evaluate them on this dataset. Instead, we compare our method with the methods consuming normalized points, including NDT-Transformer [44], PPT-Net [45], SVT-Net [42], and TransLoc3D [43]. For our method, we generate BEV images of size 200$\times$200. Following the previous works [14, 37], we train our method only with the RobotCar training dataset and test the method on the test set. For all the compared methods, we directly use the results from their papers. Tab. VI shows that our BEVPlace++ outperforms other methods including the transformer-based ones with large margins. In particular, our method generalizes well to U.S, R.A, and B.D subsets, while other methods have relatively large performance degradation.
@@ VI-B Loop Closing
Similar to the setup in place recognition, we test the methods with models trained on sequence “00” of KITTI. We perform evaluation on every single sequence. For each frame in a sequence, we perform place retrieval in the former frames with the nearest 100 frames excluded.
@@
Tab. VII presents a quantitative comparison of average precision, F1 max score, max recall at 100% precision, mean translation errors, and mean rotation errors. BEVPlace++ achieves an average precision over 90% on both datasets. While LCDNet and Logg3DNet also show high precision on KITTI, their precision and F1 max scores are significantly lower on NCLT. Conversely, BoW3D fails to generalize to NCLT. Notably, BEVPlace++ achieves a significantly higher maximum recall at 100% precision on sequence ”08” compared to other methods. This high level of precision is crucial because false positive detections can introduce irreversible errors into downstream tasks such as loop correction and map updating. Ensuring a high recall at perfect precision means that BEVPlace++ can reliably recognize places without mistakenly identifying incorrect matches, thus maintaining the integrity and accuracy of subsequent processing steps in autonomous navigation systems. Additionally, BEVPlace++ demonstrates low mean translation and rotation errors on both datasets, specifically below 0.5 meters and 1.5 degrees, respectively. These small errors are significant because they indicate a high level of initial accuracy in the localization process. The full PR curves are illustrated in Fig. 8. As shown, BEVPlace++ outperforms all baseline methods in terms of PR curve evaluation. Especially on NCLT, BEVPlace++ has a much higher curve.
TABLE VIII: Complete Global Localization Performance on different datasets.
Dataset Sequence BoW3D LCDNet BEVPlace++
Recall (%) $\uparrow$ SR(%) $\uparrow$ $\bar{e}_{t}$ (m) $\downarrow$ $\bar{e}_{r}$ () $\downarrow$ Recall (%) $\uparrow$ SR(%) $\uparrow$ $\bar{e}_{t}$ (m) $\downarrow$ $\bar{e}_{r}$ () $\downarrow$ Recall (%) $\uparrow$ SR(%) $\uparrow$ $\bar{e}_{t}$ (m) $\downarrow$ $\bar{e}_{r}$ () $\downarrow$
NCLT 2012-02-04 14.9 3.8 1.11 2.08 60.5 58.5 0.37 1.15 95.3 95.6 0.32 1.06
2012-03-17 10.7 3.0 1.02 2.44 54.2 52.0 0.37 1.26 94.2 95.1 0.33 1.18
2012-06-15 6.5 1.1 1.23 2.62 44.2 40.0 0.49 1.28 90.2 90.9 0.42 1.11
2012-09-28 5.0 0.6 0.92 1.98 34.9 32.2 0.44 1.27 88.9 89.8 0.46 1.23
2012-11-16 5.2 0.3 1.27 2.36 31.7 28.8 0.47 1.54 91.3 90.2 0.44 1.65
2013-02-23 7.5 1.1 1.05 2.14 10.9 6.8 0.50 1.62 87.8 88.5 0.37 1.05
MCD ntu day_02 0.0 0.0 - - 45.6 29.9 1.06 0.91 83.1 77.9 0.75 1.08
day_10 0.0 0.0 - - 53.5 37.3 0.88 0.90 90.2 88.8 0.61 1.03
night_04 0.0 0.0 - - 50.3 30.3 1.01 0.98 86.6 80.9 0.62 1.01
night_08 0.0 0.0 - - 46.2 33.7 0.89 0.94 88.9 85.7 0.64 1.08
night_13 0.0 0.0 - - 48.9 36.1 0.92 0.99 86.4 85.9 0.72 1.21
VI-C Complete Global Localization
@@
Here, we conduct experiments to evaluate the accuracy of complete global localization, which estimates 3-DoF poses against a pre-built map without knowing the initial pose. For each query, we retrieve its Top-1 match from the database via place recognition and then compute the global pose using pose estimation. Tab. VIII shows the recall of place retrieval, localization success rates, mean translation error, and mean rotation error. Our BEVPlace++ generalizes better on NCLT and MCD compared to LCDNet and BoW3D, demonstrating its superiority across different sensor configurations and diverse environments. Notably, the localization success rate of BEVPlace++ on NCLT may be higher than the retrieval recall rate. This occurs when BEVPlace++ successfully estimates the pose of a query even when the distance between the query and the Top-1 match is larger than 5 meters. It should be noted that BEVPlace++ does not use pose supervision, making it much easier for deployments than methods like LCDNet [20].
@@ VI-D Runtime Analysis
We compare the runtime of the methods on a desktop equipped with an RTX 3090 GPU and an Intel quad-core 3.40 GHz i5-7500 CPU. Tab. IX shows the running time of each stage on the KITTI dataset. Our BEVPlace++ comprises simple residual and NetVLAD blocks, enabling it to run fast and achieve an average frequency of over 40 Hz for place recognition. For complete global localization, it takes about 41.6 ms for pairwise feature extraction, 0.24 ms for place retrieval, and 12.7 ms for pose estimation, achieving an average frequency of over 20 Hz. Considering that the frequency of LiDAR scans is usually set to 10 Hz, BEVPlace++ can operate in real time.
TABLE IX: Recall at Top-1 on the KITTI dataset.
Feature Extract. (ms) Place Retrieval (ms) Pose Estim. (ms)
@@
In addition to running in real-time, BEVPlace++ is also lightweight. The BEVPlace++ model is only 17 MB, significantly smaller than the model size of LCDNet 138 MB. Moreover, the BEV image averages about 20.4 KB per frame, which substantially reduces memory consumption. In contrast, LCDNet stores raw point cloud data, requiring approximately 4.0 MB per frame on KITTI.
Figure 9: Performance of A-LOAM with the LCDNet (left) compared to BEVPlace++ (right) on sequence 08 of the KITTI dataset. (a), (b) are detected loops (green lines) under 100% precision. x indicate the start point of the trajectory. (c), (d) show the absolute translation errors of LCDNet and BEVPlace++, respectively. (e), (f) are the optimized point cloud map by two methods at the same crossroads.

### v3

@@ V-B Evaluation Metrics
We use different metrics to evaluate different tasks. For place recognition, we use the recall at top-1 following [26, 77, 14]. For each query, we find its Top-1 match from the database. According to a distance threshold of 5 meters [26], we determine whether the prediction is a true positive (TP), a false negative (FN), or a false positive (FP). The recall rate is defined as the ratio of TP over the actual positives, i.e.
@@
For loop closing with loop closure detection and pose estimation, we use the Precision-Recall (PR) curve, average precision, F1 max score, max recall at 100% precision, mean translation errors, and mean rotation errors. The PR curve shows both recall rates and precision, where precision is the ratio of true positives (TP) over all predicted positives, i.e.
@@
For global localization, we compute the localization success rate (SR) under a threshold of $(2m,5^{\circ})$ following [18]. All the successful localizations are used to compute the mean translation errors and mean rotation errors.
VI Experiments
TABLE II: Recall at Top-1 on the KITTI dataset.
@@ VI Experiments
TABLE III: Recall at Top-1 on the rotated KITTI dataset.
@@
In this section, we conduct experiments to evaluate the performance of our method in terms of place recognition, loop closing, and global localization. We compare our method with state-of-the-art methods including M2DP [29], CVTNet [22], Logg3D-Net [24], BoW3d [52], RING++ [57], LCDNet [17], LCRNet [18], and EgoNN [16]. Among them, BoW3d, RING++, LCDNet, LCRNet, and LCRNet can estimate poses, while the other 3 methods can only perform place recognition and loop closure detection.
Implementation details. For all baseline methods, we reproduce their results using their open-source implementations with default setups [^2] https://github.com/LiHeUA/M2DP,https://github.com/BIT-MJY/CVTNet,https://github.com/YungeCui/BoW3D,https://github.com/csiro-robotics/LoGG3D-Net,https://github.com/robot-learning-freiburg/LCDNet,https://github.com/nubot-nudt/LCR-Net,https://github.com/jac99/Egonn.. For BEVPlace++, we use ResNet34 as the backbone in REM. The ResNet is cropped to retain the first three layers (up to conv3_x), resulting in an output channel number $C$ of 128. The number of rotations $N_{R}$ in REM is empirically set to 8. The point cloud crop range $D$ is set to 40 meters, and the grid size $g$ for BEV image generation is 0.4 meters. Consequently, the BEV image has a size of $200\times 200$. We train BEVPlace++ with the AdamW optimizer for 50 epochs. The learning rate is initially set as 1e-4 and decays by a factor of 0.5 every 10 epochs. The weight decay is set to 1e-3. The method is trained on an RTX 3090 GPU.
VI-A Place Recognition
Performance on KITTI. We only train the methods on KITTI dataset using the database of sequence “00”, containing 3000 frames. We apply data augmentation to all methods by randomly rotating the point clouds to improve the robustness to view changes. As can be seen in Tab. II, our BEVPlace++ outperforms M2DP, Scan Context, BoW3D, and CVTNet with large margins. Logg3d-Net, LCDNet, LCRNet, and EgoNN achieve comparable performance to BEVPlace++. However, they perform much worse than BEVPlace++ on the challenging sequence “08” which has a large number of challenging reverse loops.
Robustness to view changes. To test the robustness against rotational changes, we randomly rotate all the query and database point clouds around the z-axis with an angle range of $[0,2\pi)$ to simulate view changes. As shown in Tab. III, BEVPlace++ maintains the highest recall rates, benefiting from our rotational invariant global descriptor designs. CVTNet, Logg3d-Net, LCDNet, and EgoNN are also robust to rotations to some extent. However, BoW3D’s performance significantly degenerated compared to those without rotation changes.
TABLE IV: Generalization Performance of Recall at Top-1 on NCLT and MCD datasets.
@@ VI-A Place Recognition
TABLE V: Recall rates on the benchmark dataset.
@@
TABLE VI: Average precision and F1 max score of loop closure on KITTI and NCLT. † 3-DoF methods.
@@
RING++ [57] 0.809 0.776 15.9 1.12 1.01 0.852 0.799 16.4 1.02 1.27 0.843 0.801 25.3 0.97 0.84 0.836 0.808 6.0 1.29 1.35
LCDNet [36] 0.952 0.918 12.2 0.21 0.47 0.633 0.342 0.0 0.39 1.20 0.621 0.362 0.0 0.37 1.16 0.684 0.321 0.0 0.37 1.26
LCRNet [18] 0.680 0.674 3.0 1.41 1.08 0.823 0.736 2.0 1.87 1.56 0.828 0.763 6.4 1.29 1.37 0.847 0.770 2.7 1.88 1.41
EgoNN [16] 0.539 0.689 31.1 0.38 0.69 0.154 0.285 1.4 0.82 1.70 0.091 0.216 0.3 1.54 1.93 0.105 0.222 0.3 1.58 1.20
BEVPlace++† 0.999 0.984 76.4 0.35 0.57 0.963 0.912 24.9 0.34 1.09 0.969 0.916 34.5 0.36 1.19 0.935 0.859 31.2 0.40 1.17
@@
LCRNet [18] 0.839 0.805 2.7 1.77 1.92 0.613 0.548 4.6 1.70 1.82 0.132 0.258 0.0 1.95 2.46 0.482 0.603 0.0 1.79 2.04
EgoNN [16] 0.115 0.243 0.5 1.37 1.99 0.216 0.321 0.8 1.79 1.83 0.036 0.067 0.0 1.76 2.23 0.010 0.101 0.0 1.63 2.30
BEVPlace++† 0.955 0.901 63.4 0.40 1.19 0.957 0.894 45.3 0.40 1.61 0.839 0.733 15.8 0.40 1.10 0.959 0.887 45.5 0.44 1.05
Generalization ability and Long-term performance. We evaluate the methods on NCLT and MCD datasets using models trained on KITTI. For NCLT, we construct the database with the sequence “2012-01-15” that covers most areas of the campus. We then perform place recognition using the point clouds of other sequences, including the one collected in 2013 across a year. For MCD, we build the database with the sequence “ntu_day_01” and perform place recognition using other sequences, including the three night sequences. These two datasets are collected using different types of LiDAR scanners compared to those used in KITTI and their point clouds are sparser. Tab. IV shows the recall rates at Top-1 on the two datasets. As can be seen, BEVPlace++ achieves high recalls on NCLT regardless of season changes. The compared methods rather have much lower recall rates. BEVPlace++ consistently outperforms other methods on MCD with day-night changes.
@@
Performance on sparse point maps of RobotCar and Inhouse datasets. The two datasets provide point clouds downsampled to 4096 points and normalized to range $[-1,1]$. They contain coarse position ground truth. While the compared methods need raw points input or accurate pose supervision, we do not evaluate them on this dataset. Instead, we compare our method with the methods using normalized points, including NDT-Transformer [40], PPT-Net [41], SVT-Net [38], and TransLoc3D [39]. For our method, we generate BEV images of size 200$\times$200. Following the previous works [14, 33], we train our method only with the RobotCar training dataset and test the method on the test set. For all compared methods, we directly use the results from their papers. Tab. V shows that our BEVPlace++ outperforms other methods including the transformer-based ones with large margins. In particular, our method generalizes well to U.S, R.A, and B.D subsets, while other methods have relatively large performance degradation.
TABLE VII: Global localization results on Apollo, Mulran, NCLT, and MCD datasets. † 3-DoF methods.
Apollo Mulran NCLT 2012-02-02 NCLT 2012-03-17
Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$
BoW3D[52] 0.0 0.0 - - 3.6 1.8 0.57 0.46 14.9 3.8 1.11 2.08 10.7 3.0 1.02 2.44
RING++ [57] 83.2 76.4 0.62 0.44 79.1 65.7 0.71 0.49 60.0 56.8 1.11 1.23 51.3 42.2 1.29 1.33
LCDNet[17] 51.1 53.4 0.56 1.01 35.8 11.2 1.09 1.29 60.5 58.5 0.37 1.15 54.2 52.0 0.37 1.26
LCRNet[18] 67.4 91.8 0.11 1.57 57.9 39.1 0.99 1.57 56.2 48.6 1.05 1.57 85.6 65.6 0.73 1.58
EgoNN[16] 87.9 93.9 0.22 0.47 71.6 76.0 0.41 0.70 46.3 13.0 1.09 2.19 44.6 13.7 1.05 2.19
BEVPlace++ 88.5 81.6 0.55 0.29 83.1 80.7 0.49 0.40 95.3 95.6 0.32 1.06 94.2 95.1 0.33 1.18
NCLT 2012-06-15 NCLT 2012-09-28 NCLT 2012-11-16 NCLT 2013-02-23
Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$
BoW3D [52] 6.5 1.1 1.23 2.62 10.7 3.0 1.02 2.44 5.2 0.3 1.27 2.36 7.5 1.1 1.05 2.14
RING++ [57] 52.1 40.3 1.25 1.68 31.2 30.3 1.32 1.75 35.5 28.7 1.29 1.71 26.9 19.1 1.33 1.75
LCDNet [17] 44.2 40.0 0.49 1.28 34.9 32.2 0.44 1.27 31.7 28.8 0.47 1.54 10.9 6.8 0.50 1.62
LCRNet[18] 43.2 34.8 0.52 1.59 33.9 6.9 0.91 1.59 43.2 14.2 0.93 1.58 37.1 18.2 1.01 1.54
EgoNN[16] 39.2 11.7 1.11 2.13 39.5 12.0 1.11 2.16 33.5 8.3 1.12 2.35 32.0 8.3 1.05 2.18
BEVPlace++† 90.2 90.9 0.42 1.11 88.9 89.8 0.46 1.23 91.3 90.2 0.44 1.65 87.8 88.5 0.37 1.05
MCD ntu_day_02 MCD ntu_day_10 MCD ntu_night_04 MCD ntu_night_08
Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$ Recall (%) SR (%) $\hat{e}_{t}$ (m) $\hat{e}_{r}$ $(^{\circ})$
BoW3D [52] 0.0 0.0 - - 0.0 0.0 - - 0.0 0.0 - - 0.0 0.0 - -
RING++ [57] 42.8 40.1 1.22 1.31 49.5 42.4 1.13 1.37 44.1 40.2 1.34 1.50 52.2 49.3 1.24 1.33
LCDNet [17] 45.6 29.9 1.06 0.91 53.5 37.3 0.88 0.90 50.3 30.3 1.01 0.98 46.2 33.7 0.89 0.94
LCRNet[18] 81.5 76.3 0.41 1.57 75.7 70.8 0.50 1.57 50.3 45.0 0.44 1.57 63.2 58.4 0.50 1.57
EgoNN[16] 60.2 18.4 1.12 2.39 59.1 27.0 1.07 2.17 30.0 10.6 0.99 2.33 60.0 27.3 1.04 2.22
BEVPlace++† 83.1 77.9 0.75 1.08 90.2 88.8 0.61 1.03 86.6 80.9 0.62 1.01 88.9 85.7 0.64 1.08
@@ VI-B Loop Closing
Similar to the setup in place recognition, we test the methods with models trained on sequence “00” of KITTI. We perform evaluation on every single sequence. For each frame in a sequence, we perform place recognition in the former frames with the nearest 100 frames excluded [43, 20].
@@
Tab. VI presents a quantitative comparison of average precision, F1 max score, max recall at 100% precision, mean translation errors, and mean rotation errors. BEVPlace++ achieves an average precision of over 90% on both datasets. While LCDNet and Logg3DNet also show high precision on KITTI, their precision and F1 max scores are significantly lower on NCLT. LCRNet and EgoNN struggle with reverse loops and have relatively lower precision on sequence “08” of KITTI. Notably, BEVPlace++ achieves a significantly higher maximum recall at 100% precision on sequence ”08” compared to other methods. This high precision is crucial because false positive detections can introduce irreversible errors into downstream tasks such as loop correction and map updating. Ensuring a high recall at high precision means that BEVPlace++ can reliably recognize places without mistakenly identifying incorrect matches, thus maintaining the integrity and accuracy of subsequent processes in autonomous navigation systems. Additionally, BEVPlace++ demonstrates low mean translation and rotation errors on both datasets, specifically below 0.5 meters and 1.5 degrees, respectively. These small errors are significant because they indicate a high level of initial accuracy in the localization process. The full PR curves are illustrated in Fig. 7. As shown, BEVPlace++ outperforms all baseline methods in terms of PR curve evaluation. Especially on NCLT, BEVPlace++ has a much higher curve.
VI-C Global Localization
@@
This experiment evaluates the accuracy of global localization, estimating 3-DoF poses against a pre-built map without knowing the initial pose. For each query, we retrieve its Top-1 match from the database via place recognition and then compute the global pose using pose estimation. Tab. VII shows the place recognition recall, localization success rates, mean translation error, and mean rotation error. Our BEVPlace++ achieves the highest recall on all sequences and generalizes better on Mulran, NCLT, and MCD compared to BoW3d, LCDNet, LCRNet, and EgoNN, demonstrating its superiority across different sensor configurations and diverse environments. Notably, the localization success rate of BEVPlace++ on NCLT can be higher than the retrieval recall rate, when BEVPlace++ successfully estimates the pose of a query even when the distance between the query and the Top-1 match is larger than 5 meters. It should be noted that BEVPlace++ does not use pose supervision, making it much easier for deployments than methods like LCDNet, LCRNet, and EgoNN.
@@ VI-D Runtime Analysis
We compare the runtime of the methods on a desktop equipped with an RTX 3090 GPU and an Intel quad-core 3.40 GHz i5-7500 CPU. Tab. VIII shows the running time of each stage on the KITTI dataset. BEVPlace++ comprises simple residual and NetVLAD blocks, achieving an average frequency of over 40 Hz for place recognition. For global localization, it takes 41.6 ms for pairwise feature extraction, 0.24 ms for place recognition, and 12.7 ms for pose estimation, achieving a frequency of over 20 Hz. As the frequency of LiDAR scans is usually set to 10 Hz, BEVPlace++ can operate in real time.
TABLE VIII: Inference Time on The KITTI dataset.
Feature Extract. (ms) Place Recognition (ms) Pose Estim. (ms)
@@
In addition to achieving real-time performance, BEVPlace++ is also highly lightweight. The BEVPlace++ model has a size of just 17 MB, which is significantly smaller compared to LCDNet (138 MB), LCRNet (96 MB), and EgoNN (19 MB). Furthermore, the BEV image format averages approximately 20.4 KB per frame, substantially reducing memory consumption. In contrast, LCDNet, LCRNet, and EgoNN store raw point cloud data, requiring around 4.0 MB per frame on the KITTI dataset.
Figure 8: Performance of A-LOAM with the LCDNet (left) compared to BEVPlace++ (right) on sequence 08 of the KITTI dataset. (a), (b) are detected loops (green lines) under 100% precision. x indicates the start point of the trajectory. (c), (d) show the absolute translation errors of LCDNet and BEVPlace++, respectively. (e), (f) are the optimized point cloud map by two methods at the same crossroads.